what is the missing coordinate that falls on the line 2x + 3y = 12 (_,-2)
Answer by fractalier(6550)

If y = -2 on this line:
2x + 3y = 12
then we can find x by plugging in for y...
2x + 3(-2) = 12
2x - 6 = 12
2x = 18
x = 9
(9, -2)